# Re-Simulation-based Self-Supervised Learning for Pre-Training Foundation Models

### v1

@@ Abstract
Self-Supervised Learning (SSL) is at the core of training modern large machine learning models, providing a scheme for learning powerful representations that can be used in a variety of downstream tasks. However, SSL strategies must be adapted to the type of training data and downstream tasks required. We propose RS3L, a novel simulation-based SSL strategy that employs a method of re-simulation to drive data augmentation for contrastive learning. By intervening in the middle of the simulation process and re-running simulation components downstream of the intervention, we generate multiple realizations of an event, thus producing a set of augmentations covering all physics-driven variations available in the simulator. Using experiments from high-energy physics, we explore how this strategy may enable the development of a foundation model; we show how R3SL pre-training enables powerful performance in downstream tasks such as discrimination of a variety of objects and uncertainty mitigation. In addition to our results, we make the RS3L dataset publicly available for further studies on how to improve SSL strategies.
1 Introduction
Self-Supervised Learning (SSL) has been a key driver in recent machine learning (ML) advancements; through the use of data augmentation, SSL algorithms rely on pseudo labels to create representations of the data that can be highly useful and efficiently adapted for downstream tasks. Accordingly, the resulting SSL trained model can act as a powerful foundation model [1], defined here as a model that is ``pre-trained'' on a generic task, in this case one defined with SSL, and capable of being used in a variety of downstream tasks. Foundation models have quickly arisen as a strategy for describing complex tasks, especially in computer vision (e.g., [2]), natural language processing (e.g., [3]), and speech recognition (e.g., [4]). More recently, foundation models for machine learning in science have become an important and active area of research (for instance in biology [5], chemistry [6, 7, 8, 9], differential equation solving [10, 11], and cosmology [12, 13])
@@ 1 Introduction
To that end, we propose a simulation-based strategy to obtain an augmented data set that is as domain complete as possible and only limited by the knowledge encapsulated in a simulator. More specifically, in settings with stochastic simulators, we propose a strategy whereby we intervene in the middle of a simulation process, fix the upstream generated latent state of an event, and re-simulate downstream components to sample augmentations from the set of all plausible observations for a given latent state. From a programmatic standpoint, the first step fixes all initial conditions and stochastic latent variables in the simulation up to the intervention point, while the second step re-samples the stochastic latent variables in the simulation after the intervention point and thus generates a new stochastic output of the simulation conditioned on the fixed latent state at the intervention point. We dub our new strategy RS3L ([”rIz9l], ``Re-simulation-based self-supervised representation learning''). We will show that this method allows for powerful contrastive pre-training that, through altering the simulator settings for re-simulation, also aids in the mitigation of uncertainties arising from domain shift between potentially imperfect simulations and real data.
@@
In particular, we will focus on learning representations of jets, which are highly energetic, collimated streams of secondary particles. Jets are core objects in HEP; they are produced from the showering and hadronization of high-energy quarks and gluons or through the decay of heavy elementary particles like the Higgs boson, the top quark, or heavy weak gauge bosons. Jets are featured in many theories for new physics beyond the Standard Model. Hence, the ability to ``tag'' a jet, i.e., to correctly and robustly classify the particle which initiated the jet, is crucial for the success of the physics program of experiments like ATLAS [16] and CMS [17] at the Large Hadron Collider (LHC) [18] and future high energy collider experiments. Our choice of re-simulating the parton shower is motivated by a desire to characterize the progenitor of a jet, while capturing the large variability that results for each parton shower. That is, the re-simulated shower defines an augmentation to the existing shower that we claim captures all the physics that we desire downstream.
By sampling from our high-fidelity simulator to obtain augmented versions of jets, our work presents an evolution of the methods developed in JetCLR [19], which focused on empirical methods of augmentation generation, whereas our work develops a domain-complete augmentation set based concretely on the physics in the simulation. As a result, we outline a strategy through augmentations within SSL to pre-train a latent space that captures the most salient, physically motivated, features of a simulation. Within high energy physics, other SSL methods have also been explored for foundation model pre-training. For instance, the masked modeling strategies developed in BERT [20] and BEiT [21] have been adapted for set-type data in order to mask and predict particles in a jet [22], and to mask and predict particle-type information [23]. Supervised pre-training and fine-tuning strategies have also been explored for high energy physics contexts [24]. A similar strategy has been also pursued within astrophysics [25].
@@
  • The creation of a large open dataset for common development in the community [^2] https://doi.org/10.5281/zenodo.10633815.
@@ 2.1 The RS3L backbone
The above set captures both 1) implicit uncertainty in the definition of a quark or gluon parton shower, which results in a cascade of particles that can vary in number of particles and distribution of energy across the particles, and 2) the variations within parton shower simulations that are known to cover the observed variations in the data. The augmentations themselves can be interpreted physically as a construction of the parton ``wave function'' before its collapse into a measured state. The resulting space thus aims to project away the imperfect modeling of the parton shower process and of the statistical variability in our simulator, including quantum effects and ill-defined hidden parameters.
@@
Given a nominal jet, we randomly sample one of its augmentations from the above set to build a positive pair. All other possible jet pairs in the sample are considered negative pairs. We then process each jet with the backbone, which we define using graph-based architectures with built-in graph convolutions and message-passing. Graph neural networks (GNNs) have seen large success over other ML algorithms like dense or convolutional neural networks in the analysis of LHC data due to the point-cloud nature of the particles that comprise a jet or an event (for an extensive review, see [29]). Finally, each jet is embedded into an $N$-dimensional latent space. We use $N=8$ dimensions for the latent space as a balance between expressiveness and computational complexity, although higher-dimensional spaces are possible [^3] We observed that networks trained with 128 and 512 output dimensions performed similarly in the downstream tasks.. The 8D jet representations are then used in the contrastive loss function based on SimCLR. The loss function aims at aligning the 8D vectors of jets building a positive pair, and pushes apart jets from negative pairs. A temperature parameter $\tau$ regulates the relative importance of these two simultaneous training objectives and is set to 0.1. All training details are provided in Appendix A.
2.2 Understanding the contrastive space
@@
Of the eight dimensions, the most discriminating dimension between Higgs bosons jets and quarks and gluons is shown on the left of Figure 2. As a qualitative statement on the robustness of the space we derive through RS3L, we also show in the same figure an equivalent plot for the variable $N_{2}$, which is a theory-motivated function of the energy flow of particles in the jet [30]. The latter represents a ratio between the compatibilities of the jet to have one or, respectively, two prongs, which are regions of collimated radiation. As a powerful jet tagging variable, $N_{2}$ is commonly used in searches for processes involving hadronically decaying Higgs bosons [31, 32, 33] or (additional) weak gauge bosons [34, 35] to reduce the overwhelming background where jets arise from quantum chromodynamics (QCD) interactions. In both plots, the distributions of Higgs jets are shown in magenta and the distributions of QCD jets (jets initiated from quarks and gluons) are shown in other colors (see also Appendix A). The main (upper) panel shows the RS3L or, respectively, $N_{2}$ feature for the nominal parton shower scenario, while the bottom panels show the ratios of the varied distributions to the nominal distribution.
Figure 2: (Left) One of eight features derived in the RS3L pre-training. (Right) Jet substructure variable $N_{2}$.
@@
In these plots, smaller ratios indicate smaller differences in the learned representation for jets with different simulator configurations, and therefore smaller systematic uncertainties arising from the parton showering model chosen. Large trends are observed in these ratio when comparing different simulator configurations in the case of $N_{2}$, while the trends are mitigated in the RS3L space thanks to the employed self-supervision strategy of aligning nominal jets with their augmented versions. The one-dimensional distributions for all RS3L features are included in Appendix A.
@@
Figure 3: Corner plots for the eight outputs of RS3L, split up into Higgs boson (reds) and QCD jets (blues). Only small correlations are observed among feature pairs.
@@
Figure 4: 2D visualization of the 8D RS3L space, derived via t-SNE dimensionality reduction. Top: A good class separation is seen between Higgs jets and QCD (quark and gluon) jets. Bottom left: Jets shown by parton shower model for Pythia8 and Herwig7 for a RS3L space trained with Herwig7 augmentations. Bottom left: The same for a RS3L space trained without Herwig7 augmentations. The congruence of the different parton shower models is visibly worse in the right-hand scenario.
@@
Finally, we visually probe the entire jet representation contained in the 8D RS3L space by performing a reduction to two dimensions using the t-SNE algorithm [37]. The plot on the top of Figure 4 shows this 2D space for gluons, quarks, and Higgs jets. Again, strong separation between Higgs and QCD jets as well as strong clustering for a given class is visible [^4] Notably, the model tries to push apart jets from two different Higgs bosons. However, as all Higgs bosons share common characteristics, this objective is hard to meet, leading to the observed clustering.. On the bottom left, we show the space divided into jets showered with Pythia8 and, respectively, Herwig7. At the population level, it is clear that different showering configurations occupy similar regions in the RS3L space and, hence, our approach provides decent robustness against domain shift. The bottom right of Figure 4 shows the t-SNE 2D space of a RS3L training that did not include Herwig7 augmentations. Compared to the RS3L space obtained with all augmentations, this space has visibly smaller overlap (congruence) between Pythia8-showered and Herwig7-showered jets, speaking to larger uncertainties incurred by domain shift.
2.3 Fine-tuning on top of the RS3L backbone
@@
We now study the use of RS3L as a backbone for in-distribution and out-of-distribution classification tasks, and we quantify performance and uncertainties. For all fine-tunings, we consider a sample comprising the same types of augmentations used for training the respective RS3L space. We then compare the RS3L pre-training + fine-tuning performance to the performance of a fully-supervised network with the same architecture. For a fair comparison, we also train the fully-supervised network on a sample comprising the same types of augmentations as were used to train the RS3L space. The fully-supervised network, therefore, learns to marginalize over the augmentations, not explicitly exploiting the information that the nominal and augmented jet come from the same initial particle [^5] We point out that the standard approach for fully supervised algorithms at experiments such as ATLAS and CMS is to train on samples comprising only jets from a nominal parton shower configuration..
2.3.1 In-distribution classification and robustness

### v2

# Re-Simulation-based Self-Supervised Learning for Pre-Training Physics Foundation Models

@@ Abstract
Self-Supervised Learning (SSL) is at the core of training modern large machine learning models, providing a scheme for learning powerful representations that can be used in a variety of downstream tasks. However, SSL strategies must be adapted to the type of training data and downstream tasks required. We propose RS3L (“Re-simulation-based self-supervised representation learning”), a novel simulation-based SSL strategy that employs a method of re-simulation to drive data augmentation for contrastive learning in the physical sciences, particularly, in fields that rely on stochastic simulators. By intervening in the middle of the simulation process and re-running simulation components downstream of the intervention, we generate multiple realizations of an event, thus producing a set of augmentations covering all physics-driven variations available in the simulator. Using experiments from high-energy physics, we explore how this strategy may enable the development of a foundation model; we show how RS3L pre-training enables powerful performance in downstream tasks such as discrimination of a variety of objects and uncertainty mitigation. In addition to our results, we make the RS3L dataset publicly available for further studies on how to improve SSL strategies.
I Introduction
Self-Supervised Learning (SSL) has been a key driver in recent machine learning (ML) advancements; through the use of data augmentation, SSL algorithms rely on pseudo labels to create representations of the data that can be highly useful and efficiently adapted for downstream tasks. Accordingly, the resulting SSL trained model can act as a powerful foundation model [1], defined here as a model that is “pre-trained” on a generic task, in this case one defined with SSL, and capable of being fine-tuned for a variety of purposes. Foundation models have quickly arisen as a strategy for describing complex tasks, especially in computer vision (e.g., [2]), natural language processing (e.g., [3]), and speech recognition (e.g., [4]). More recently, foundation models for machine learning in science have become an important and active area of research, for instance in biology [5], chemistry [6, 7, 8, 9], differential equation solving [10, 11], cosmology [12, 13], and, finally, high-energy physics [14], not all of which, however are necessarily SSL-based.
@@ I Introduction
[…]
@@
In particular, we will focus on learning representations of jets, which are highly energetic, collimated streams of secondary particles. Jets are core objects in HEP; they are produced from the showering and hadronization of high-energy quarks and gluons or through the decay of heavy elementary particles like the Higgs boson, the top quark, or heavy weak gauge bosons. Jets are featured in many theories for new physics beyond the Standard Model. Hence, the ability to “tag” a jet, i.e., to correctly and robustly classify the particle which initiated the jet, is crucial for the success of the physics program of experiments like ATLAS [17] and CMS [18] at the Large Hadron Collider (LHC) [19] and future high energy collider experiments. For instance, one of the canonical classification tasks in HEP and a prerequisite for the precise understanding of Higgs boson couplings to bottom quarks [20, 21, 22, 23, 24] is the discrimination between jets originating from Higgs bosons and jets originating from QCD. Our choice of re-simulating the parton shower is motivated by a desire to characterize the progenitor of a jet, while capturing the large variability that results for each parton shower. That is, the re-simulated shower defines an augmentation to the existing shower that we claim captures all the physics that we desire downstream.
By sampling from our high-fidelity simulator to obtain augmented versions of jets, our work presents an evolution of the methods developed in JetCLR [25] and [26], which focused on empirical methods of augmentation generation (such as rotations), whereas our work develops a domain-complete augmentation set based concretely on the physics in the simulation. As a result, we outline a strategy through augmentations within SSL to pre-train a latent space that captures the most salient, physically motivated, features of a simulation. Within high energy physics, other SSL methods have also been explored for foundation model pre-training. For instance, the masked modeling strategies developed in BERT [27] and BEiT [28] have been adapted for set-type data in order to mask and predict particles in a jet [29], and to mask and predict particle-type information [30]. Supervised pre-training and fine-tuning strategies have also been explored for high energy physics contexts [31]. A similar strategy has been also pursued within astrophysics [32].
@@
  • The creation of a large open dataset for common development in the community https://doi.org/10.5281/zenodo.10633815.
@@ II.1 The RS3L backbone
The above set captures both 1) implicit uncertainty in the definition of a quark or gluon parton shower, which results in a cascade of particles that can vary in number of particles and distribution of energy across the particles, and 2) the variations within parton shower simulations that are known to cover the observed variations in the data. The augmentations themselves can be interpreted physically as a construction of the parton “wave function” before its collapse into a measured state. The resulting space thus aims to “project away”, i.e., disregard, the imperfect modeling of the parton shower process and the statistical variability in our simulator, including quantum effects and ill-defined hidden parameters.
@@
The RS3L space is trained with 5M events using a 50%/50% admixture of QCD jets and Higgs jets. Given a nominal jet, we randomly sample one of its augmentations to build a positive pair. All other possible jet pairs in a minibatch, which consists of 100 nominal and 100 augmented jets, are considered negative pairs. We then process each jet with the backbone, which we define using graph-based architectures with built-in graph convolutions and message-passing. Graph neural networks (GNNs) have seen large success over other ML algorithms like dense or convolutional neural networks in the analysis of LHC data due to the point-cloud nature of the particles that comprise a jet or an event (for an extensive review, see [36]). Finally, each jet is embedded into an $N$-dimensional latent space. We use $N=8$ dimensions for the latent space as a balance between expressiveness and computational complexity, although higher-dimensional spaces are possible [^1] We observed that networks trained with 128 and 512 output dimensions performed similarly in the downstream tasks without notable improvement over $N=8$.. The 8D jet representations are then used in the contrastive loss function based on SimCLR. The loss function aims at aligning the 8D vectors of jets building a positive pair, and pushes apart jets from negative pairs. A temperature parameter $\tau$ regulates the relative importance of these two simultaneous training objectives and is set to 0.1.
II.2 Data augmentations and training input
Hard-scattering events for $\text{p}\text{p}\to\text{Z}+\text{jet}$ and $\text{p}\text{p}\to\text{H}\text{Z}$, with $\text{H}\to\text{b}\bar{\text{b}}$ and $\text{Z}\to\nu\bar{\nu}$, from which we sample the jets for training the networks, are generated at a center-of-mass energy of $\sqrt{s}=13\,\text{TeV}$ using MadGraph5_aMC@NLO v3.4.0 [37] with the NNPDF30_nlo_nf_5_pdfas parton distribution function [38] at leading-order accuracy in QCD. The j and, respectively, the Higgs boson are generated within a $p_{\mathrm{T}}$ range of 400-600 GeV and within |$\absolutevalue{\eta}<0.1$. The hard-scattering events are then passed through a parton shower simulation and are reconstructed using the Delphes v3.4.3pre1 detector simulation [39] with a CMS detector-like geometry. For gluons and the Higgs, the first splitting or, respectively, decay is generated at matrix-element level with a minimum $\Delta R=\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}>0.001$. All further splittings are generated by the subsequent parton shower.
II.3 Network architecture
$h_{\mathrm{embed}}=\mathrm{MLP}_{\mathrm{embed}}(X)$
$h_{\mathrm{DEC1}}=\textsc{DynamicEdgeConv}(h_{\mathrm{embed}}|k=24)$
$h_{\mathrm{DEC2}}=\textsc{DynamicEdgeConv}(h_{\mathrm{DEC1}}|k=24)$
$h_{\mathrm{DEC3}}=\textsc{DynamicEdgeConv}(h_{\mathrm{DEC2}}|k=24)$
$h_{\mathrm{enc}}=\mathrm{MLP}_{\mathrm{enc}}(h_{\mathrm{DEC3}})$
$\bm{z}=\textsc{GlobalSumPool}(h_{\mathrm{enc}})\ .$ (1)
The input feature set $X$ is given by 15 features per particle, listed in Tab. 1. These input features are not normalized over a minibatch, but their magnitudes are comparable among the features and $\mathcal{O}(1)$, thereby regularizing the network and leading to a similar effect as a pre-transformation that enforces a mean of 0 and standard deviation of 1.
For embedding the 15 particle features into a higher-dimensional latent space, two fully-connected layers with dimensions (128/128) are used in $\textsc{MLP}_{\text{embed}}$. After a series of three convolutions with $k=24$, $\mathrm{MLP}_{\mathrm{enc}}$ consists of four fully-connected layers with dimensions (64/32/32/8). Outputs of all neurons in both MLPs are activated with an exponential linear unit [44]. The final building block, GlobalSumPool, aggregates the eight features of each particle in the jet from the last layer of $\mathrm{MLP}_{\mathrm{enc}}$ to obtain the final eight features $\bm{z}$ characterizing the jet in the RS3L space.
Table 1: Description of the particle input features.
Particle feature Description
$\log{p_{\text{T}}}$ ) logarithm of the transverse momentum
$\log{E}$ ) logarithm of the energy
$\log{(E/E^{\text{Jet}})}$ ) logarithm of $E$ normalized w.r.t. the jet
$\Delta\eta$ difference in pseudorapidity w.r.t. the jet
$\Delta\phi$ difference in $\phi$ w.r.t. the jet
$\Delta R$ distance in $\eta$-$\phi$ plane from the jet axis
$\tanh d_{0}$ hyperbolic tangent of the transverse impact parameter
$\tanh d_{z}$ hyperbolic tangent of the longitudinal impact parameter
$q$ particle charge
isPhoton 1 if particle is a photon, else 0
isMuon 1 if particle is a muon, else 0
isElectron 1 if particle is an electron, else 0
isCH 1 if particle is a charged hadron, else 0
isNH 1 if particle is a neutral hadron, else 0
The network will try to push apart dissimilar jets (negative pairs) on the $N-1$-dimensional hypersphere with unit radius given by the $N$ input features, while trying to align similar jets (positive pairs) and characterize them by similar features. The competition between these two effects is regulated by a temperature parameter $\tau$. We studied this parameter based on the performance obtained when using the RS3L space to distinguish Higgs from QCD jets. Specifically, we trained RS3L spaces with $\tau$ ranging from $0.05$ to $0.5$ and fine-tuned these networks on the Higgs vs. QCD classification task with a fixed backbone. A fixed backbone offers the best insight into the performance of the RS3L pre-training step because it projects jets into a one-dimensional classification space using only the information derived during the self-supervised step. We observed that the Higgs vs. QCD binary classification performance was optimal with $\tau=0.1$.
II.4 Fine-tuning and fully-supervised trainings
The MLP placed on top of the RS3L architecture is given by one additional fully-connected layer with 8 input dimensions and a single sigmoid activation function at the output. The backbone network architecture used in the fine-tuning is the exact same as the fully-supervised. The only difference between the approaches is that in the fine-tuning method the base graph trained with RS3L is loaded as a hot-start for the classification and the MLP is added on top of the backbone network. We consider two configurations of fine-tuning the RS3L space: a “fixed” graph (where only the MLP parameters are floating) and a “floating” graph (where all the MLP and graph parameters are floating). The admixture of QCD/H/W, as well as augmentations composition, are always consistent between the RS3L fine-tuning and full supervision. The RS3L space and classification steps are trained on independent datasets.
III Results
III.1 Understanding the contrastive space
@@
Of the eight dimensions, the most discriminating dimension between Higgs bosons jets and quarks and gluons is shown on the left of Figure 3. As a qualitative statement on the robustness of the space we derive through RS3L, we also show in the same figure an equivalent plot for the variable $N_{2}$, which is a theory-motivated function of the energy flow of particles in the jet [45]. The latter represents a ratio between the compatibilities of the jet to have one or, respectively, two prongs, which are regions of collimated radiation. As a powerful jet tagging variable, $N_{2}$ is commonly used in searches for processes involving hadronically decaying Higgs bosons [46, 47, 48] or (additional) weak gauge bosons [49, 50] to reduce the overwhelming background where jets arise from quantum chromodynamics (QCD) interactions. In both plots, the distributions of Higgs jets are shown in magenta and the distributions of QCD jets (jets initiated from quarks and gluons) are shown in other colors (see also section II). In particular, QCD jets are divided by their flavor content and whether the initial parton was a single quark or a gluon. The main (upper) panel shows the RS3L or, respectively, $N_{2}$ feature for the nominal parton shower scenario, while the bottom panels show the ratios of the varied distributions to the nominal distribution.
Figure 3: (Left) One of eight features derived in the RS3L pre-training. (Right) Jet substructure variable $N_{2}$. The main (upper) panel shows the distributions for the nominal parton shower scenario. The ratio panels show the difference between the respective varied distributions and the nominal distribution. For FSR, the up and down variations form a band around the nominal distribution.
@@
In these plots, ratios closer to 1 indicate smaller differences in the learned representation for jets with different simulator configurations, and therefore smaller systematic uncertainties arising from the parton showering model chosen. Large trends are observed in these ratio when comparing different simulator configurations in the case of $N_{2}$, while the trends are mitigated in the RS3L space thanks to the employed self-supervision strategy of aligning nominal jets with their augmented versions.
@@
Figure 4: Corner plots for the eight outputs of RS3L, split up into Higgs boson (reds) and QCD jets (blues). Only small correlations are observed among feature pairs, as indicated by the Pearson correlation coefficients provided in each subplot.
@@
Figure 5: 2D visualization of the 8D RS3L space, derived via t-SNE dimensionality reduction. Top: A good class separation is seen between Higgs jets and QCD (quark and gluon) jets. Bottom left: Jets shown by parton shower model for Pythia8 and Herwig7 for a RS3L space trained with Herwig7 augmentations. Bottom right: The same for a RS3L space trained without Herwig7 augmentations. The congruence of the different parton shower models is visibly worse in the right-hand scenario.
@@
Finally, we visually probe the entire jet representation contained in the 8D RS3L space by performing a reduction to two dimensions using the t-SNE algorithm [52]. The plot on the top of Figure 5 shows this 2D space for gluons, quarks, and Higgs jets. Again, strong separation between Higgs and QCD jets as well as strong clustering for a given class is visible. Notably, the model tries to push apart jets from two different Higgs bosons. However, as all Higgs bosons share common characteristics, this objective is hard to meet, leading to the observed clustering. On the bottom left, we show the space divided into jets showered with Pythia8 and, respectively, Herwig7. At the population level, it is clear that different showering configurations occupy similar regions in the RS3L space and, hence, our approach provides decent robustness against domain shift. The bottom right of Figure 5 shows the t-SNE 2D space of a RS3L training that did not include Herwig7 augmentations. Compared to the RS3L space obtained with all augmentations, this space has visibly smaller overlap (congruence) between Pythia8-showered and Herwig7-showered jets, speaking to larger uncertainties incurred by domain shift.
III.2 Fine-tuning on top of the RS3L backbone
@@
We now study the use of RS3L as a backbone for in-distribution and out-of-distribution classification tasks, and we quantify performance and uncertainties. For all fine-tunings, we consider a sample comprising the same types of augmentations used for training the respective RS3L space. We then compare the RS3L pre-training + fine-tuning performance to the performance of a fully-supervised network with the same architecture. For a fair comparison, we also train the fully-supervised network on a sample comprising the same types of augmentations as were used to train the RS3L space. The fully-supervised network, therefore, learns to marginalize over the augmentations, not explicitly exploiting the information that the nominal and augmented jet come from the same initial particle. We point out that the standard approach for fully supervised algorithms at experiments such as ATLAS and CMS is to train on samples comprising only jets from a nominal parton shower configuration.
III.2.1 In-distribution classification and robustness